Input the equation for the following line in standard form Ax + By = C. Reduce all fractional answers to lowest terms.
A line with y-intercept of 2 that is parallel to 2x + y = -5.

Answer by mananth(16946)

2x + y = -5.
slope o f this line = -2
ithe required line will have same slope
passing through (0,2)
y-2=-2(x-0)
y-2=-2x
y+2x=2